What is the percent ionization in a 0.300 M solution of formic acid (HCOOH) (Ka 1.78 x 10-4)?
Answer #1

2H2(g) + O2(g) = 2420 (g) (Acid) = 0.300M ka = 1.78X10 4 dissociated concentration =( = 6.300 X 1-76x157* = 0.0073 M Initial